### v1

# User-customizable Shared Control for Fine Teleoperation via Virtual Reality

Abstract
Shared control can ease and enhance a human operator’s ability to teleoperate robots, particularly for intricate tasks demanding fine control over multiple degrees of freedom. However, the arbitration process dictating how much autonomous assistance to administer in shared control can confuse novice operators and impede their understanding of the robot’s behavior. To overcome these adverse side-effects, we propose a novel formulation of shared control that enables operators to tailor the arbitration to their unique capabilities and preferences. Unlike prior approaches to “customizable” shared control where users could indirectly modify the latent parameters of the arbitration function by issuing a feedback command, we instead make these parameters observable and directly editable via a virtual reality (VR) interface. We present our user-customizable shared control method for a teleoperation task in $\mathrm{SE}(3)$, known as the buzz wire game. A user study is conducted with participants teleoperating a robotic arm in VR to complete the game. The experiment spanned two weeks per subject to investigate longitudinal trends. Our findings reveal that users allowed to interactively tune the arbitration parameters across trials generalize well to adaptations in the task, exhibiting improvements in precision and fluency over direct teleoperation and conventional shared control.
I Introduction
Teleoperation is the foundation behind many robotic applications. These applications range from preventing human presence in hazardous environments to providing healthcare services through remote patient care and robot-assisted surgery. Nevertheless, remotely controlling a complex robot warrants a certain level of operator skill and domain expertise in order to successfully complete challenging teleoperation tasks [1]. Without this baseline, a novice operator may be exposed to detrimental levels of physical and cognitive workload [2], potentially resulting in catastrophic task failures [3]. A common means of relieving this burden on the operator is to employ an autonomous controller that continuously assists the human user by sharing control over the robot [4].
While shared control alleviates excess workload exerted on operators, it may also give rise to new issues that impact task performance and impede dexterity training [5]. One prominent example of such an issue is model misalignment. This phenomenon occurs when the division in control between the human and autonomy creates a misunderstanding in how the operator expects the robot to behave [6]. In direct teleoperation, model misalignment may arise whenever a robot has higher degrees-of-freedom (DoF) than the user’s control interface [7]. Despite this interface asymmetry in DoF, a user with adequate training will still gradually build a mental model of the robot’s behavior. In contrast, the extra layer of arbitration introduced in shared control exacerbates the user’s difficulty in comprehending the system. A communication medium for user feedback is therefore essential to how an operator understands the arbitration [4].
Aside from equipping operators with suitable feedback, another integral aspect of shared control is user customizability. Over a long-term interaction, operators will experience variations in their capabilities, preferences, and environments [8]. Shared control must account for these changes and adapt accordingly, especially in terms of how control authority is assigned. Prior works have developed promising frameworks for adaptive shared control by enabling users to either interactively customize the parameters that characterize the arbitration function [9] or suggest corrections for the resulting robot behavior [10, 11]. However, these works rely entirely on the robot’s legible state to convey the effects of a user’s modifications to the arbitration procedure.
Figure 1: Illustration of the Kinova Gen 3 robotic arm being teleoperated to play the buzz wire game by a user wearing a virtual reality (VR) headset. The VR controller held by the operator will be represented as a handle in the virtual world, where its simulated effects also have a direct consequence on the physical counterpart played by the robot.
In this paper, we establish user-customizable shared control by directly communicating the internal parameters of the arbitration function to an operator for refinement. By unveiling these parameters, users can better understand the arbitration process and thus effectively edit its inner workings for personalized outcomes. We ground this idea in a teleoperation task, where operators must remotely control a 7-DoF robotic arm in virtual reality (VR) to complete the popular buzz wire game (see Fig. 1). The buzz wire game is a suitable testbed due to the need for precise control over both translational and rotational robot motion. This game also requires significant eye-hand coordination and sustained operator focus [12, 13]. Furthermore, the game is frequently utilized for skill training in domains that would benefit from robot teleoperation, such as surgery [14]. To assess the quality and longevity of the proposed shared control method, we conducted a user study that explores patterns in participant performance over repeated interactions.
The key contributions of this paper are as follows:
  • A novel mathematical framework that factors in user feedback to formalize user-customizable shared control;
  • A comprehensive demonstration of this framework within the context of a teleoperation task in ${\mathrm{SE}(3)}$, involving a real 7-DoF robotic arm and a VR interface as the bidirectional communication channel;
  • Results from a longitudinal user study with 12 subjects teleoperating the robotic arm to perform the buzz wire game over four sessions spread across two weeks.
II Related Work
Sensory feedback can be supplied across multiple modalities in teleoperation. For example, haptic devices are widely used in bilateral teleoperation to simultaneously enhance an operator’s task efficacy and encourage growth in skill [4, 2]. Other modalities, e.g., audio, force, and visual feedback, have enjoyed similar success [15, 16, 17]. In recent years, VR headsets have proven advantageous in complex manipulation tasks where high-DoF robots are remotely controlled [18, 19, 20] . Moreover, VR interfaces offer operators the opportunity to modify the inner workings of the teleoperated robot, which is an asset we exploit in this work.
In addition to communication, shared control must also consider arbitration, i.e., how to “blend” a user’s inputs with the autonomous controller’s outputs [4]. An arbitration function is typically described by a set of parameters that determine how control is allocated. These parameters could be fixed, e.g., to primarily favor the human leading, or they could be dynamically updated. In the latter case, a vast bulk of the shared control literature resorts to heuristics for parameter updates, such as safety, fluency, agreement, and confidence [21, 22, 23]. User-personalized parameters for shared control can also be configured by balancing out the operator’s capability to independently complete a teleoperation task with their need for autonomous assistance [24, 25]. However, the user is unable to configure arbitration parameters in these approaches if no bidirectional communication is available.
Heuristics-based arbitration may be suitable for certain users, but it could also lead to reduced satisfaction or performance in others. To identify the exact form of the “optimal” arbitration function and its corresponding parameters is an intractable problem in real-world teleoperation [9]. Therefore, it may be more appropriate to empower users with the ability to customize the arbitration parameters online. While previous approaches to customizable shared control have utilized verbal feedback [9, 26] or corrections [10, 11] to grant this ability, they solely depended on the robot’s observable behavior to explain the control-sharing. Relying on the robot’s appearance and movement to communicate any underlying autonomous assistance is often insufficient [27]. Instead, our paper opts for greater transparency by externalizing the factors that influence arbitration and making them editable to operators.
III User-customizable Shared Control
In the following, we propose a mathematical framework for shared control that enables a user to dictate how the robot autonomy should provide teleoperation assistance.
III-A Formulation
The shared control system considered in this work is driven by the following interaction data. At time $t$, a human operator’s measured state, $\mathbf{x}_{h}(t)\in\mathbb{R}^{n_{h}}$ (e.g., joystick deflections), is fed into a control interface to produce control commands, $\mathbf{u}_{h}(t)\in\mathbb{R}^{m_{h}}$, for direct robot teleoperation. Similarly, a robot’s state and the autonomy’s assistive commands can be denoted as $\mathbf{x}_{r}(t)\in\mathbb{R}^{n_{r}}$ (e.g., joint angles, Cartesian poses) and $\mathbf{u}_{r}(t)\in\mathbb{R}^{m_{r}}$, respectively.
Under this setup, the output commands of a shared control system adhere to an arbitration process:
$\mathbf{u}_{sc}(t)=\bm{\beta}_{\bm{\theta}}(\mathbf{u}_{r}(t),\mathbf{u}_{h}(t)),$ (1)
where $\bm{\beta}(\cdot)$ is an arbitration function parameterized by time-varying, $\bm{\theta}(t)$, or time-invariant variables, $\bm{\theta}$. As an example, the standard linear blending scheme could be written as:
$\bm{\beta}_{\bm{\theta}}(\mathbf{u}_{r}(t),\mathbf{u}_{h}(t))=(1-\alpha_{\bm{\theta}})\mathbf{u}_{r}(t)+\alpha_{\bm{\theta}}\mathbf{u}_{h}(t),$ (2)
where $\alpha_{\bm{\theta}}\in[0,1]$ is the blending variable and is also a function parameterized by $\bm{\theta}$ [9]. Note that $\alpha$ is frequently treated as a scalar, but in our approach we will adopt a matrix representation to facilitate more granular authority over the system behavior. Let $\mathbf{A}_{\bm{\theta}}=\text{diag}(\alpha_{1},\ldots,\alpha_{m_{r}})$ be a positive definite diagonal arbitration matrix, with $\forall i,\alpha_{i}\in[0,1]$, such that the blending scheme is now:
$\bm{\beta}_{\bm{\theta}}(\mathbf{u}_{r}(t),\mathbf{u}_{h}(t))=(\mathbf{I}-\mathbf{A}_{\bm{\theta}})\mathbf{u}_{r}(t)+\mathbf{A}_{\bm{\theta}}\mathbf{u}_{h}(t),$ (3)
where $\mathbf{I}$ is an identity matrix.
Selecting the parameters, $\bm{\theta}$, of an arbitration function, $\bm{\beta}_{\bm{\theta}}(\cdot)$, is critical in shaping the robot’s assistive behavior. Instead of manually selecting a fixed choice of $\bm{\theta}$, an “optimal” selection is often derived by solving an optimization problem with an objective function, $\Gamma$, such as:
$\Gamma=\mathcal{L}_{T}(\mathbf{x}_{r}(t))+\sum_{t=0}^{T-1}\mathcal{L}\left(\mathbf{x}_{r}(t),\bm{\beta}_{\bm{\theta}}\left(\mathbf{u}_{r}(t),\mathbf{u}_{h}(t)\right)\right),$ (4)
where $\mathcal{L}_{T}(\cdot)$ is a terminal cost for a horizon, $T$ (e.g., error between the intended and actual target robot states), and $\mathcal{L}(\cdot)$ is an incremental cost (e.g., human effort). Finding a time-invariant $\bm{\theta}$ to minimize $\Gamma$ will ignore the dynamic nature of the human-robot collaboration. Even if a solution is found with time-varying $\bm{\theta}(t)$, the core issue remains that $\Gamma$ might not be the true cost function of the shared control, $\Gamma^{*}$, nor would a tractable solution to $\Gamma^{*}$ necessarily exist [9].
III-B Feedback-informed User Optimization
Figure 2: Diagram of the user-customizable shared control architecture.
To circumvent the intractable problem of identifying $\Gamma^{*}$ and its “optimal” parameters, we present a method that allows users to personalize the arbitration function by configuring their preferred $\bm{\theta}$. The key insight behind our formulation is to establish a bidirectional communication channel between the human and robot. We formalize this channel as a vector-valued function $\bm{\psi}(\cdot)$ (dropping $t$ for brevity):
$\bm{\theta}\leftarrow\bm{\psi}(\mathbf{v}_{r},{v}_{h}),$ (5)
where the robot administers feedback, $\mathbf{v}_{r}$, to the operator about the shared control, while the user performs parameter updates, $\Delta\bm{\theta}$, via an interface action, $v_{h}$.
For instance, a scalar blending quantity, $\alpha$, could be derived from the following piece-wise function [9, 25]:
$\alpha$ =$=\begin{cases}0&C\leq\theta_{1},\\ \frac{\theta_{3}}{\theta_{2}-\theta_{1}}\cdot C&\theta_{1}<C\leq\theta_{2},\\ \theta_{3}&C>\theta_{2},\end{cases}$ (6)
$\text{where}\quad\theta_{j}$ $=\psi_{j}(\mathbf{v}_{r},v_{h}),\quad j=1:K.$
The $\delta\theta_{j}$ in the $K$ parameter mappings, $\psi_{j}(\mathbf{v}_{r},v_{h})=\theta_{j}\pm\delta\theta_{j}$, could be obtained from discrete verbal commands made by the operator, $v_{h}\in\mathbb{Z}^{+}$ (e.g., “more/less assistance”). Here, the autonomy’s confidence in its estimation of a user’s goal, $C$, is independent of the resulting arbitration level. We highlight that most previous efforts at user customization of $\theta_{j}$ have solely depended on the robot’s legible state, $\mathbf{v}_{r}=\mathbf{x}_{r}$, for operator feedback [9, 26].
In the multi-dimensional arbitration scenario, one could choose to represent $\mathbf{A}_{\bm{\theta}}$ in a more general form:
$\mathbf{A}_{\bm{\theta}}$ $=\operatorname{diag}(\bm{\alpha})=\operatorname{diag}(\alpha_{1},\ldots,\alpha_{m_{r}}),$ (7)
$\text{where}\quad\alpha_{i}$ $=\sum_{j=1}^{K}w_{j}\psi_{j}(\mathbf{v}_{r},v_{h}),\quad i=1,\ldots,m_{r}.$
Each diagonal element, $\alpha_{i}$, is governed by potentially more than one factor, $\theta_{j}$, resultant from $\psi_{j}(\mathbf{v}_{r},v_{h})$, as determined by the selection weight, $w_{j}$. Whether the arbitration is expressed as Eq. 6 or Eq. 7, the aspiration behind permitting operator customization of $\bm{\theta}$ is to converge on a user-elected optimum, without concretely specifying $\Gamma^{*}$. The overall framework architecture is shown in Fig. 2.
As the feedback variables, $\mathbf{v}_{r}$ and $v_{h}$, dramatically impact how users customize the shared control behavior, we suggest two improvements over prior works in how to produce these variables. First, the robot’s communication flow should illuminate more than its external state, $\mathbf{x}_{r}$. In Section IV, we demonstrate how a set of human-interpretable signals, $\mathcal{V}$ (i.e., visualizations in VR), can be incorporated into $\mathbf{v}_{r}=\{\mathbf{x}_{r},\mathcal{V}\}$ to help operators tune $\bm{\theta}$. Second, an operator’s parameter updates, $\Delta\bm{\theta}$, should have a proportionate effect on robot behavior. This proportionality can be achieved by letting the operator submit actions, $v_{h}\in[0,1]$ (e.g., using a graphical slider in VR), over the same constrained range as the arbitration parameters, $\forall j,\theta_{j}\in[0,1]$.
IV Proposed Framework Instantiation
To illustrate the operational principles of the proposed framework, this section provides an instantiation for assistive teleoperation in the scope of the buzz wire game.
IV-A Testbed: Buzz Wire Game
In the buzz wire game, the operator’s goal is to carefully guide a loop handle in $\mathrm{SE}(3)$ from start to end while minimizing collisions with the wire. The game is a well-known assessment environment for fine motor skill and eye-hand coordination [12, 13]. Moreover, 3D buzz wire games are advantageous when evaluating dexterity tasks that benefit from stereoscopic depth perception [12]. As precise motor control, eye-hand coordination, and 3D spatial awareness are critical capabilities for teleoperation in general, we deem this game to be a pertinent testbed. Immersive VR versions of the game have also been validated as an effective simulation for operator training, e.g., in stroke rehabilitation [28, 29]. Hence, we believe a VR setup of the 3D game serves as a suitable experimental arena for our work.
Fig. 1 demonstrates our buzz wire game replica for robot teleoperation. The layout is comprised of a Kinova Gen 3 arm with a loop handle extension at its end-effector to complete the actual game, as well as a person wearing an HTC Vive headset to perform the same task in VR with a handheld controller. A model-based perspective of the buzz wire game and the robot are also constructed in the VR scene. An electric circuit board imitated the game’s “buzz” feature by signaling collisions to the operator in VR. Multimodal feedback cues were triggered upon collision, with red lights flashing and vibrations in the handheld input device.
IV-B Shared Assistance in Teleoperation
In teleoperation, a common strategy is to control the robot’s end-effector pose, $\mathbf{x}_{r}\in\mathbb{R}^{6}$, by sending desired twist commands, $\dot{\mathbf{x}}_{d}\in\mathbb{R}^{6}$, via admittance control:
$\dot{\mathbf{x}}_{d}=\mathbf{M}^{-1}\int(\mathbf{k}\odot\mathbf{e}+\mathbf{d}\odot\dot{\mathbf{e}}+\mathbf{f}_{\mathrm{ext}})\,\mathrm{d}t,$ (8)
where $\mathbf{M}\in\mathbb{R}^{6\times 6}$, $\mathbf{k}\in\mathbb{R}^{6}$, and $\mathbf{d}\in\mathbb{R}^{6}$ are the pre-defined mass, stiffness, and damping terms respectively, while $\mathbf{f}_{\mathrm{ext}}\in\mathbb{R}^{6}$ is the externally applied force, $\mathbf{e}\in\mathbb{R}^{6}$ is the pose error vector, and $\odot$ denotes component-wise multiplication. In our buzz wire setup, the pose error is the difference between the handheld VR controller’s pose, $\mathbf{x}_{h}\in\mathbb{R}^{6}$, and the robot’s state, ${\mathbf{x}_{r}\in\mathbb{R}^{6}}$.
The robot autonomy can then be integrated into Eq. 8 as an assistive force. By reformulating Eq. 8 as the linear blending scheme shown in Eq. 3, we establish an admittance control policy for shared control as follows:
$\mathbf{u}_{sc}=\dot{\mathbf{x}}_{d}$ $=(\mathbf{I}-\mathbf{A}_{\bm{\theta}})\mathbf{u}_{r}+\mathbf{A}_{\bm{\theta}}\mathbf{u}_{h},$ (9)
$\mathbf{u}_{h}$ $=\mathbf{M}^{-1}\int(\mathbf{k}\odot\mathbf{e}+\mathbf{d}\odot\dot{\mathbf{e}})\,\mathrm{d}t,$ (10)
$\mathbf{u}_{r}$ $=\mathbf{M}^{-1}\int\mathbf{w}_{a}\,\mathrm{d}t.$ (11)
In this formulation, $\mathbf{f}_{\mathrm{ext}}$ has been replaced by a wrench vector, $\mathbf{w}_{a}\in\mathbb{R}^{6}$, to represent the autonomous assistance in both translational and rotational space.
IV-C Real-time Assistive Wrench
(a) Attractive force $\mathbf{f}_{i,k}$ from attractive points on the wire.
(b) Repulsive torque $\bm{\tau}_{i,k}$ from sample collision points on the wire.
Figure 3: Illustration of how the attractive and repulsive forces are computed given a control point $\mathbf{p}_{i}$ (dark blue) and environment point $\mathbf{p}_{k}$ (red).
Generating real-time autonomous assistance is vital for teleoperation to ensure a reactive user experience and smooth robot motion. To fulfil this requirement, we efficiently compute the assistive wrench, $\mathbf{w}_{a}$, using potential fields, where targets produce attractive vectors and obstacles yield repulsive vectors. Attractive and repulsive vectors are typically calculated between a pair of control and environment points [30]. In our setting, we define eight control points as the vertices of an octagon that encloses the circular end-effector. The environment points are acquired in real-time given point cloud data detected using two depth cameras.
Environment points within a pre-defined range of the control points compose a neighborhood set, $\mathcal{C}=\mathcal{C}_{\mathrm{att}}\sqcup\mathcal{C}_{\mathrm{rep}}$, where $|\mathcal{C}|=N$. The $\mathcal{C}_{\mathrm{att}}$ subset contains attractive points, whereas $\mathcal{C}_{\mathrm{rep}}$ contains repulsive points. The potential forces between an environment point, $\mathbf{p}_{k}\in{\mathcal{C}}$, and a control point, $\mathbf{p}_{i}\in\mathbb{R}^{3}$, are computed based on a logistic function:
$\mathbf{f}_{i,k}=\frac{f_{\mathrm{max}}}{1+\exp\left[-\lambda_{i,k}\left(\frac{\|\mathbf{d}\|}{\rho}-d_{i,k}\right)\right]}\ \hat{\mathbf{d}}_{i,k},$ (12)
where each function-shaping parameter ${\lambda_{i,k}\in\{\lambda_{\mathrm{rep}},\lambda_{\mathrm{att}}\}}$, ${d_{i,k}\in\{d_{\mathrm{rep}},d_{\mathrm{att}}\}}$, ${\hat{\mathbf{d}}_{i,k}=\pm\frac{\mathbf{d}_{i,k}}{\|\mathbf{d}_{i,k}\|}}$ has two possible values depending on the relationship between control point $i$ and environment point $k$ (repulsive or attractive). Hence, forces $\mathbf{f}_{i,k}$ can either be repulsive or attractive, with $f_{\mathrm{max}}$ determining the maximum magnitude of the generated force and $\rho$ adjusting the change rate in force given distance.
As the buzz wire game requires 6-DoF teleoperation, we compute the assistive torque $\bm{\tau}_{i,k}$ for rotational motion as:
$\bm{\tau}_{i,k}=\mathbf{f}_{i,k}\times(\mathbf{p}_{\mathrm{com}}-\mathbf{p}_{i})$ (13)
where $\mathbf{p}_{\mathrm{com}}$ is the center of mass of the robot’s end-effector. Fig. 3 illustrates an example of how attractive forces and repulsive torques are calculated for the buzz wire game given a pair of control and environment points $(\mathbf{p}_{i},\mathbf{p}_{k})$.
The resultant assistance wrench, $\mathbf{w}_{a}=\begin{bmatrix}\mathbf{f}_{\mathrm{net}}^{\mathsf{T}}&\bm{\tau}_{\mathrm{net}}^{\mathsf{T}}\end{bmatrix}^{\mathsf{T}}$, is obtained by aggregating a force component, ${\mathbf{f}_{\mathrm{net}}\in\mathbb{R}^{3}}$, and a torque component, $\bm{\tau}_{\mathrm{net}}\in\mathbb{R}^{3}$, each calculated as:
$\mathbf{f}_{\mathrm{net}}$ $=\sum_{i}\frac{\mathbf{f}_{i}}{\|\mathbf{f}_{i}\|}\cdot\max_{k}{\|\mathbf{f}_{i,k}\|},$ (14)
$\bm{\tau}_{\mathrm{net}}$ $=\sum_{i}\frac{\bm{\tau}_{i}}{\|\bm{\tau}_{i}\|}\cdot\max_{k}{\|\bm{\tau}_{i,k}\|},$ (15)
where ${\mathbf{f}_{i}=\sum_{k}\mathbf{f}_{i,k}}$, ${\bm{\tau}_{i}=\sum_{k}\bm{\tau}_{i,k}}$, $i\in\{1,\ldots,8\}$, and $k\in\{1,\ldots,N\}$.
IV-D User-customizable Arbitration
Figure 4: The customizable spider chart interface in virtual reality. The operator can view and modify assistance across five factors, before confirming their configuration. Each factor is scaled from $[0.1,1]$ to $[10,100]$ with a discrete minimal update step of $5$.
A VR interface acts as the basis for bidirectional communication, $\bm{\psi}(\mathbf{v}_{r},v_{h})$, in our user-customizable shared control. We created a spider chart interface, depicted in Fig. 4, to augment the robot’s feedback, $\mathbf{v}_{r}$, with a variety of human-readable indicators, $\mathcal{V}$. Specifically, three factors affecting arbitration are displayed in VR: speed, depth assistance, and turnability. Speed regulates robot velocity along axes of motion relevant to the buzz wire game (e.g., horizontally along the wire). Depth assistance governs the margin between the end-effector and the wire along the depth axis, with the aim of assisting line-of-sight occlusions. Turnability adjusts the rotational speed of the end-effector. Users can thus view $\mathbf{v}_{r}$, as well as modify the parameters, $\bm{\theta}$, by supplying edits, $v_{h}$, per factor via increment/decrement buttons.
Following the formulation in Eq. 7, the arbitration vector, $\bm{\alpha}\in\mathbb{R}^{6}$, is expressed as a linear combination:
$\bm{\alpha}=\mathbf{W}\bm{\theta}$ $=\mathbf{W}\bm{\psi}(\mathbf{v}_{r},v_{h})=\mathbf{W}\begin{bmatrix}v_{h}^{\mathrm{speed}}\\ 1-v_{h}^{\mathrm{depth}}\\ v_{h}^{\mathrm{turn}}\end{bmatrix},$ (19)
$\text{where}\quad\mathbf{W}$ $=\begin{bmatrix}1&0&1&0&0.4&0.2\\ 0&1&0&0.2&0&0\\ 0&0&0&0.8&0.6&0.8\end{bmatrix}^{\mathsf{T}}.$ (23)
The coefficient matrix, $\mathbf{W}\in\mathbb{R}^{6\times 3}$, is structured in a manner where each editable facet, $v_{h}$, directly maps to a $\theta$ with a semantically similar effect on arbitration. The weighting of elements was chosen empirically to fit the task. For example, $v_{h}^{\mathrm{turn}}$ is masked by $\mathbf{W}$ to contribute to an $80\%$ change in arbitration on rotation about the z-axis, while $v_{h}^{\mathrm{speed}}$ contributes to the remaining $20\%$. Here, the term turnability intuitively represents more influence over robot rotation than translational speed. The weighting criteria for depth assistance uses the complement of $v_{h}^{\mathrm{depth}}$, as “assistance” implies more contribution from the autonomy.
In addition to task-specific parameters $\bm{\theta}$, we expose two task-independent parameters: safety and responsiveness. These parameters adjust $f_{\mathrm{max}}$ in Eq. 12 and $\mathbf{k}$ in Eq. 10, respectively, akin to a variable impedance controller. Note that each configurable parameter is actually constrained in the range $[0.1,1]$ to prevent a “full autonomy” configuration.
V Experiment
To gauge the effectiveness of the proposed framework in enhancing teleoperation performance, we conducted a between-subjects study involving 12 subjects (3 female; aged 20-34). Subjects were asked to complete the buzz wire game by teleoperating a robotic arm in VR. All participants provided written consent prior to data collection.
V-A Experiment Protocol
Participants were allocated into one of three groups, each designed to evaluate a different control strategy: (1) direct teleoperation (teleop), which is implemented by setting $\mathbf{A}_{\bm{\theta}}\,{=}\,\mathbf{I}$ in Eq. 9; (2) Shared control with a heuristics-based arbitration function (sc), which is described in Section V-B; and (3) our user-customizable shared control method (sc_user). Participants were unaware of the distinctions between each group. It is also worth noting that half of the 12 subjects reported no background in robotics. This equal distribution persisted across the control strategies.
Before commencing the experiment, participants in groups sc and sc_user were provided with identical written explanations of the five arbitration factors, described in Section IV-D. After each trial, these participants could review the updated arbitration factors for the subsequent trial through the VR interface (Fig. 4). Only sc_user subjects could edit these factors. All subjects were instructed to complete the game as safely and quickly as possible.
Each group consisted of four participants, and every subject carried out five consecutive trials of the buzz wire game in a single session using their assigned control strategy. This procedure was repeated for a total of four sessions. To evaluate the learning rates of operators over an extended period, a minimum of one workday separated each session, resulting in a two-week duration for the entire experiment. Additionally, the last session involved a distinct wire testbed, as portrayed in Fig. 5. This alteration aimed to assess whether any teleoperation skills gained are transferable to a related, yet different task. Subjects also answered post-session surveys on their user experience.
Figure 5: New wire testbed for the last session to assess skill generalization.
V-B Heuristics-based Shared Control
We implemented a heuristics-based assist-as-needed controller as the reference method, sc, for our experiments. The arbitration parameters, $\bm{\theta}$, in this controller are iteratively updated according to the operator’s performance after each trial. We linearly update the parameters with a change rate akin to the multi-trial update function from [24]:
$\bm{\theta}_{t+1}=(1+\bm{\chi}_{t})\bm{\theta}_{t},$ (24)
where $t$ is the trial index and $\bm{\chi}_{t}\in\mathbb{R}^{3}$ is the change rate vector. $\bm{\chi}_{t}$ is updated using a non-linear function based on the two most recent trials:
$\chi_{t}^{i}=\chi_{\mathrm{nom}}\frac{\bar{r}_{t}^{i}-r_{d}^{i}}{r_{d}^{i}}\left(\frac{|\bar{r}_{t}^{i}-r_{d}^{i}|}{|\bar{r}_{t-1}^{i}-r_{d}^{i}|}\right)^{\displaystyle\operatorname{sign}(r_{d}^{i}-\bar{r}_{t}^{i})},$ (25)
where $r_{d}^{i}$ is the $i^{\mathrm{th}}$ component of the desired error vector $\mathbf{r}_{d}$, ${\bar{r}^{i}_{t}}$ is the $i^{\mathrm{th}}$ component of the average error vector ${\bar{\mathbf{r}}_{t}}$ during trial $t$, and $\chi_{\mathrm{nom}}$ is the nominal change rate, which determines the maximum change rate within one update step.
Distance from the end-effector to the closest point along the wire is a suitable performance metric to denote safety in the buzz wire game. This error is conveniently contained in the assistive wrench, $\mathbf{w}_{a}$, such that $\bar{\mathbf{r}}_{t}$ can be expressed as:
$\bar{\mathbf{r}}_{t}=\mathbf{W}^{\dagger}\overline{\mathbf{w}}_{a,t},$ (26)
where $\overline{\mathbf{w}}_{a,t}\in\mathbb{R}^{6}$ is the average wrench from trial $t$, $\mathbf{W}^{\dagger}$ is the pseudo-inverse of the coefficient matrix in Eq. 19, and $\mathbf{r}_{d}$ is pre-calculated from an expert demonstration.
V-C Evaluation Metrics
Performance evaluation for the buzz wire game primarily revolves around two measures: successful task completion time and number of collisions. A trial is deemed successful and collisions are counted when the participant teleoperates the robot’s end-effector to the game end without encountering a fatal failure (when collision forces exceed a threshold). Aside from task time and collisions, we also examine a user’s jerk in handling the VR controller (2nd derivative of input velocity commands) to capture “smoothness” [31]. Additionally, we explore the learning trajectories of users, i.e., each user’s change in “skill” metrics relative to their first session. Lastly, we investigate a user’s perceptions of the different control modes in terms of skill, success, and difficulty.
Therefore, the hypotheses for this study are as follows:
  • H1: Subjects will report higher ratings on skill, success, and difficulty for sc_user than teleop and sc.
  • H2: There will be faster task success times and less collisions for sc_user than teleop and sc.
  • H3: Participants in sc_user will exhibit better teleoperation performance relative to their ‘baseline’ session when compared against users from teleop and sc.
VI Results
In the following, we report on our results and test our three hypotheses. When evaluating the learning trajectories of participants, we term the first session as the ‘baseline’, the combined second and third sessions as ‘training’, and the last session as ‘transfer’. Unless otherwise stated, mixed ANOVAs were performed to analyze the effects of control modes (between-subjects) and sessions (within-subjects) on the relevant evaluation metrics. The ‘baseline’ session was left out of the ANOVAs, leaving only the combined ‘training’ sessions to compare against the ‘transfer’ condition.
VI-A Subjective Results
TABLE I: Subjective results across sessions (1 = “very low” or “failure”, 10 = “very high” or “perfect”)
Sessions Mode Skill $\uparrow$ Success $\uparrow$ Difficulty $\downarrow$
1 (baseline) sc 7.5 $\pm$ 1.0 7.5 $\pm$ 1.3 4.8 $\pm$ 2.2
sc_user 6.3 $\pm$ 0.5 6.0 $\pm$ 1.8 6.5 $\pm$ 1.7
teleop 6.8 $\pm$ 1.0 7.5 $\pm$ 1.0 5.5 $\pm$ 2.4
2 & 3 (training) sc 6.1 $\pm$ 3.9 6.0 $\pm$ 1.6 5.9 $\pm$ 1.8
sc_user 7.1 $\pm$ 3.8 7.4 $\pm$ 1.4 6.1 $\pm$ 1.4
teleop 7.6 $\pm$ 3.6 7.5 $\pm$ 1.4 5.8 $\pm$ 2.4
4 (transfer) sc 6.0 $\pm$ 0.8 6.8 $\pm$ 1.0 5.0 $\pm$ 1.8
sc_user 7.3 $\pm$ 0.5 7.0 $\pm$ 0.8 5.8 $\pm$ 2.1
teleop 6.5 $\pm$ 2.4 6.5 $\pm$ 2.4 7.0 $\pm$ 3.6
TABLE II: Subject responses to overall experiment questions (1 = “strongly disagree”, 5 = “strongly agree”).
Question Teleop SC SC-User
I found the spider chart helpful - 3.5 $\pm$ 1.0 4.3 $\pm$ 1.0
Most would learn this robot quickly 4.0 $\pm$ 0.8 3.8 $\pm$ 0.5 4.5 $\pm$ 0.6
The last session was more difficult 3.3 $\pm$ 1.7 2.5 $\pm$ 0.6 3.0 $\pm$ 1.2
Subject ratings to post-session questions on skill (“How would you rate your skill?”), success (“How successful were you?”), and difficulty (“How hard did you have to work?”) are shown in Table I. For the ‘baseline’ session, sc subjects self-reported better ratings than the other two groups, especially in terms of perceived difficulty, with a mean difference of $1.7$ between sc and sc_user. There is a tendency toward more favorable responses for teleop and sc_user in the ‘training’ sessions, with teleop receiving the best ratings across all categories. Nevertheless, sc_user has the superior ‘transfer’ according to perceived skill ($7.3\pm 0.5$) and success ($7.0\pm 0.8$). The teleop users found the new wire testbed most difficult to complete ($7.0\pm 3.6$), with a mean score difference of $2.0$ from sc and $1.2$ from sc_user. Despite these patterns, no main or interaction effects were revealed from the mixed ANOVAs performed on each question category. Based on these tests, we cannot substantiate H1.
Survey questions on the entire experiment are summarized in Table II. Indeed, the last session is voted as marginally more difficult by teleop participants. Subjects in the sc_user group also viewed operating the robot as “quicker” to grasp than the other modes. Interestingly, sc_user participants considered the spider chart to be more “helpful” ($4.3\pm 1.0$) than the sc chart ($3.5\pm 1.0$), which had autonomously computed values and could not be edited. All sc_user individuals “found the VR interface easy to edit” ($5.0$, not in table).
VI-B Quantitative Results
(a) Time-to-success per session
(b) Time-to-success per trial
Figure 6: Time-to-success demonstrated for participants across (a) sessions and (b) trials. Gray shaded area represents the session with a new test bed.
TABLE III: Quantitative results across sessions
Sessions Mode Collisions Time (s) No. Fails
1 (baseline) sc 1.74 $\pm$ 2.6 51.4 $\pm$ 24.8 1
sc_user 2.74 $\pm$ 1.9 38.4 $\pm$ 8.5 1
teleop 2.10 $\pm$ 1.8 35.8 $\pm$ 9.8 0
2 & 3 (training) sc 1.60 $\pm$ 2.0 53.1 $\pm$ 12.8 6
sc_user 1.50 $\pm$ 1.3 35.0 $\pm$ 8.5 0
teleop 1.54 $\pm$ 1.4 29.3 $\pm$ 8.0 1
4 (transfer) sc 2.26 $\pm$ 1.9 43.2 $\pm$ 5.4 1
sc_user 1.45 $\pm$ 1.1 25.5 $\pm$ 6.3 0
teleop 1.70 $\pm$ 1.4 28.5 $\pm$ 10.3 0
Table III summarizes the quantitative results of the study. In the first ‘baseline’ session, the sc group had the least number of collisions ($1.74\pm 2.6$), whereas teleop had the fastest time-to-success ($35.8\text{s}\pm 9.8\text{s}$). On the other hand, sc_user subjects exhibited notable progress in ‘training’ sessions, as there was a major reduction in average number of collisions, yielding the lowest collision count ($1.50\pm 1.3$). The teleop group continued to maintain the fastest success times ($29.3\text{s}\pm 8.0\text{s}$). Although in the ‘transfer’ session, sc_user participants attained the swiftest completion times ($25.5\text{s}\pm 6.3\text{s}$) and the fewest collisions ($1.45\pm 1.1$).
[…]
(a) Collisions vs baseline per session
(b) Control interface jerk vs baseline per session
Figure 7: Differences in collision occurrences (a) and control interface jerk (b) relative to the subject’s average baseline performance in the first session.
[…]
VII Discussion
A few notable findings emerged from the buzz wire experiment. First, while the subjective and absolute quantitative results were comparable for sc_user and teleop during ‘training’ sessions, sc_user operators outperformed the other two control modes in measures of relative improvement. For example, sc_user subjects had consistently less collisions and produced less jerky control inputs than their ‘baseline’ session, which were statistically significant differences over teleop and sc in the ‘transfer’ condition. Indeed, sc_user obtained markedly better results across all subjective and quantitative (absolute and relative) metrics in the new game. Altogether, these results reinforce the benefits of offering users an intuitive means of customizing shared control, particularly when there are variations in the task.
Figure 8: User-customized vs heuristics-based arbitration parameters at the conclusion of the experiment.
Another key observation is the inferior performance of the sc policy. Comparing the distributions of arbitration parameters between sc_user and sc reveals higher variance in the sc_user group’s final selection of $\bm{\theta}$, as depicted in Fig. 8. The invariance in sc is due to the performance heuristic $\bar{\mathbf{r}}_{t}$ from Eq. 26 being solely dependent on the distance between the end-effector and wire. An alternative update routine based on $\bar{\mathbf{r}}_{t}$ could prioritize factors that sc_user participants gravitated toward, such as turnability and responsiveness. However, any heuristics-based method for automatically adapting $\bm{\theta}$ would still converge to a similar parameter distribution across users, without accounting for individual preferences. Shared control policies learnt from large demonstration datasets, e.g., using reinforcement learning [32], would also suffer from the lack of personalization to unseen users. Moreover, labeling arbitration parameters for a dataset requires regular human intervention and is thus difficult to simulate or acquire.
Lastly, we believe our formulation for user-customizable shared control can be applied to a broad range of teleoperation contexts. This formulation not only modeled arbitration factors that are specific to the task (depth assistance, speed, and turnability), but also incorporated task-independent factors, such as safety to tune the potential force magnitude and responsiveness to adjust the stiffness in admittance control. One could also extend the proposed shared control formulation to treat the dimensions of latent actions learnt from expert demonstrations as user-adjustable [11, 32, 33]. The main challenge here would be in communicating these dimensions through $\mathcal{V}$, and supplying interface actions, $v_{h}$, to modify their values. Each independent latent dimension may not represent an interpretable factor of variation, nor occupy a constrained space of values for easy editing via $v_{h}$. We defer investigation into this idea for future work.
VIII Conclusions
In this paper, we introduced a mathematical shared control framework for users to customize the arbitration process. We presented an instantiation of this framework for a teleoperation task in $\mathrm{SE}(3)$ and conducted a longitudinal study spanning two weeks per subject to evaluate the proposed method. Our user-customizable shared control method enhanced teleoperation outcomes, including collision frequency and input jerk, while also accommodating the baseline skill levels of different users. Furthermore, operators transferred to generalizations in the teleoperation task better when they could edit the inner workings of the shared control. We believe these preliminary results hold promise for long-term adaptability to diverse variations in users and tasks. Future research directions will increase the participant sample size, explore a wider array of teleoperation tasks, and consider more flexible representations of the arbitration parameters for communication and user editing.